# Graph Neural Network based Log Anomaly Detection and Explanation

### v1

@@ 6.9. Sensitivity Analysis
The Number of Convolutional Layers: $L$ is a potentially important parameter as it determines how many convolutional layers to use in OCDiGCN. Figure 7 (top row) depicts PR AUC and ROC AUC for the five benchmark datasets when $L$ is varied from 1 to 5. We found that $L=1$ yields consistently good performance. As the value of $L$ is increased, there is only a slight enhancement in the resulting performance or even degradation, while the associated computational burden increases substantially. We thus recommend and use $L=1$.
The Embedding Dimension $d$: From Table 7 (middle row) , one can see that $d=128$ yields good performance on Spirit, Hadoop, HDFS and Thunderbird, while further increasing $d$ obtains negligible performance improvement or even degradation. However, an increase of $d$ on BGL leads to significantly better performance. One possible reason is that BGL is a complex dataset wherein anomalies and normal instances are not easily separable on lower dimensions.
@@
We observe that there is no significant improvement in performance when using a value of $k=2$ in comparison to $k=1$. It is important to recognize that a node exhibits 0th-order proximity with itself and 1st-order proximity with its immediately connected neighbors. If $k=2$, a node can directly aggregate information from its 2-order neighbours. As described in Table 1, graphs generated from logs usually contain a limited number of nodes, varying to 6 to 34. Therefore, this is no need to utilise the Inception Block, which was originally designed to handle large graphs in (Tong et al., 2020).
6.10. Runtime Analysis
Note that traditional machine learning methods, including PCA, OCSVM, IForest and HBOS, usually perform log anomaly detection in a transductive way. In other words, they require the complete dataset beforehand and do not follow a train-and-test strategy. In contrast, neural network based methods, such as DeepLog, LogAnomaly, AutoEncoder, and Logs2Graphs, perform log anomaly detection in an inductive manner, namely following a train-and-test strategy.
@@ 6.10. Runtime Analysis
Figure 8. Runtime for all eight methods on all datasets, wherein HDFS, BGL, and Thunderbird have been downsampled to 10,000 graphs. Runetimes are averaged over 10 repetitions. We report the training time per epoch for neural network based methods.
7. Threats to Validity
@@
Limited Datasets. Our experimental protocol entails utilizing five publicly available log datasets, which have been commonly employed in prior research on log-based anomaly detection. However, it is important to acknowledge that these datasets may not fully encapsulate the entirety of log data characteristics. To address this limitation, our future work will involve conducting experiments on additional datasets, particularly those derived from industrial settings, in order to encompass a broader range of real-world scenarios.
@@
Purity of Training Data. The purity of training data is usually hard to guarantee in practical scenarios. Although Logs2Graphs is shown to be robust to very small contamination in the training data, it is critical to improve the model robustness by using techniques such as adversarial training (Bai et al., 2021) in the future.
Graph Construction. The graph construction process, especially regarding the establishment of edges and assigning edge weights, adheres to a rule based on connecting consecutive log events. However, this rule may be considered overly simplistic in certain scenarios. Therefore, more advanced techniques will be explored to construct graphs in the future.
8. Conclusions
@@
Extensive results on five benchmark datasets reveal that Logs2Graphs is at least comparable to and often outperforms state-of-the-art log anomaly detection methods such as DeepLog and LogAnomaly. Furthermore, a comparison to a similar method for graph-level anomaly detection on undirected graphs demonstrates that directed log graphs lead to better detection accuracy in practice.
Acknowledgements.

### v3

@@ 6.9 Sensitivity Analysis
The Number of Convolutional Layers: $L$ is a potentially important parameter as it determines how many convolutional layers to use in OCDiGCN. Figure 7 (top row) depicts PRC AUC and ROC AUC for the five benchmark datasets when $L$ is varied from 1 to 5. We found that $L=1$ yields consistently good performance. As the value of $L$ is increased, there is only a slight enhancement in the resulting performance or even degradation, while the associated computational burden increases substantially. We thus recommend to set $L=1$.
The Embedding Dimension $d$: From Table 7 (middle row) , one can see that $d=128$ yields good performance on Spirit, Hadoop, HDFS and Thunderbird, while further increasing $d$ obtains negligible performance improvement or even degradation. However, an increase of $d$ on BGL leads to substantially better performance. One possible reason is that BGL is a complex dataset wherein anomalies and normal instances are not easily separable on lower dimensions.
@@
We observe that there is no large improvement in performance when using a value of $k=2$ in comparison to $k=1$. It is important to recognize that a node exhibits 0th-order proximity with itself and 1st-order proximity with its immediately connected neighbors. If $k=2$, a node can directly aggregate information from its 2nd-order neighbours. As described in Table 1, graphs generated from logs usually contain a limited number of nodes, varying to 6 to 34. Therefore, there is no need to utilise the Inception Block, which was originally designed to handle large graphs in Tong et al. (2020).
6.10 Runtime Analysis
Traditional machine learning methods, including PCA, OCSVM, IForest and HBOS, usually perform log anomaly detection in a transductive way. In other words, they require the complete dataset beforehand and do not follow a train-and-test strategy. In contrast, neural network based methods, such as DeepLog, LogAnomaly, AutoEncoder, and Logs2Graphs, perform log anomaly detection in an inductive manner, namely following a train-and-test strategy.
@@ 6.10 Runtime Analysis
Figure 8: Runtime for all eight methods on all datasets, wherein HDFS, BGL, and Thunderbird have been downsampled to 10,000 graphs. Runetimes are averaged over 10 repetitions. We report the training time per epoch for neural network based methods.
7 Threats to Validity
@@
Limited Datasets. Our experimental protocol entails utilizing five publicly available log datasets, which have been commonly employed in prior research on log-based anomaly detection. However, it is important to acknowledge that these datasets may not fully encapsulate the entirety of log data characteristics. To address this limitation, our future work will conduct experiments on additional datasets, particularly those derived from industrial settings, in order to encompass a broader range of real-world scenarios.
@@
Purity of Training Data. The purity of training data is usually hard to guarantee in practical scenarios. Although Logs2Graphs is shown to be robust to very small contamination in the training data, it is critical to improve the model robustness by using techniques such as adversarial training Bai et al. (2021) in the future.
Graph Construction. The graph construction process, especially regarding the establishment of edges and assigning edge weights, adheres to a rule based on connecting consecutive log events. However, this rule may be considered overly simplistic in certain scenarios. Therefore, application-specific techniques will be explored to construct graphs in the future.
8 Conclusions
@@
Extensive results on five benchmark datasets reveal that Logs2Graphs is at least comparable to and often outperforms state-of-the-art log anomaly detection methods such as DeepLog and LogAnomaly. Furthermore, a comparison to a similar method for graph-level anomaly detection on undirected graphs demonstrates that using directed log graphs leads to better detection accuracy in practice.
CRediT authorship contribution statement
Zhong Li: Conceptualization, Methodology, Validation, Investigation, Software, Writing - original draft, Visualisation, Project Administration. Jiayang Shi: Methodology, Validation, Software, Writing - original draft. Matthijs van Leeuwen: Methodology, Validation, Writing - review & editing, Funding acquisition.
Declaration of competing interest
The author(s) declared no potential conficts of interest with respect to the research, authorship and/or publication of this article.
Data availability
For reproducibility, all code and datasets are provided on GitHub  [^3] https://github.com/ZhongLIFR/Logs2Graph.
Acknowledgements